[figure: 9.7 cm]

PHENOMENOLOGICAL THEORY OF NEW MULTIPLE SPIN

DENSITY WAVES IN FCC TRANSITION METALS

T. Uchida^{a} and Y. Kakehashi^{b} aHokkaido Inst. Tech., Sapporo 006-8585 Japan, bMax-Planck Institut für PKS, Nöthnitzer Str. 38, D-01187 Dresden

Recent molecular-dynamics calculations and the first-principles band calculations predicted a new multiple spin density wave state whose principal waves consist of 3-Q(Q = 0.6) in the cubic γ-Fe. To clarify such a possibility from the phenomenological point of view, we investigated the stability using the Ginzburg-Landau free energy. Taking into account the terms up to the 4-th order, we obtained the magnetic phase diagram in the space of the Landau parameters. Our diagram shows up that the phase change NM → AF → 3- Q̂ → 3-Q with increasing volume, which is consistent with the microscopic prediction, is possible in addition to the change NM → AF → Helical. Here NM, AF, 3-Q̂ denote the nonmagnetic, the first-kind antiferromagnetic, and the 3-Q(Q = 1) states, respectively.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : T. Uchida

Address for correspondence : Hokkaido Institute of Technology, Maeda 7-15-4-1, Teine-ku, Sapporo 006-8585, Japan Tel. +81 11 681 2161 Fax +81 11 681 3622

Email address : uchida@hit.ac.jp